OpenStudy (aaronandyson):

An area is paved with square tiles of a certain size and the number required is 128.If the tiles had been 2 cm shorter each way,200 tiles would have been needed to pave the same area .Find the size of the larger tiles.

OpenStudy (aaronandyson):

@perl

OpenStudy (aaronandyson):

@perl

OpenStudy (perl):

Let x by x be the area of each tile, before shrinking them

OpenStudy (perl):

Let x be the side of the tile, and y be the area. Area of one tile is x^2. You need 128 of them to cover the area y. 128 * x^2 = y the dimension each tile is decreased by 2 200 ( x - 2) (x-2) = y

OpenStudy (perl):

128 * x^2 = y 200 ( x - 2) (x-2) = y 200 * ( x^2 - 4x + 4) = y not substitute 200 x^2 - 800x + 800 = 128 x^2

OpenStudy (perl):

can you finish it now?

OpenStudy (perl):

@Aaron

OpenStudy (perl):

you there?